OpenStudy (anonymous):

The students of a school went on a picnic by bus. The number of students in each of x buses was the same. The number of rows of seats in each bus was 1 more than the number of buses. The number of students who sat on each row of seats was 2 more than the number of rows of seats in each bus. The expression below shows the total number of students in all the buses: x(x + 1)(x + 3) How can the total number of rows of seats in all the buses combined be calculated

OpenStudy (anonymous):

answer choices By multiplying x and (x + 1) By multiplying x and (x + 3) By adding (x + 1) and (x + 3) By adding (x + 1) and (x + 2)

OpenStudy (anonymous):

@jim_thompson5910 can you please help?

jimthompson5910 (jim_thompson5910):

how far did you get?

OpenStudy (anonymous):

I don't know where to begin.. but I read the problem and I'm thinking it's the 1st answer. @jim_thompson5910

jimthompson5910 (jim_thompson5910):

why the first one?

OpenStudy (anonymous):

There are x buses. The number of rows in a bus is 1 more than the number of buses. now i'm not so sure :/ @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yep, there are x buses

jimthompson5910 (jim_thompson5910):

how many seats does one bus have?

OpenStudy (anonymous):

This is confusing... like i'm lost @jim_thompson5910

jimthompson5910 (jim_thompson5910):

"The number of rows in a bus is 1 more than the number of buses" there are x buses so there are ______ seats in one bus

OpenStudy (anonymous):

I know it's the first answer but I can't necessarily explain it. but when you multiply x(x+1) you get x^2+1 . @jim_thompson5910

jimthompson5910 (jim_thompson5910):

we have x buses each bus has x+1 seats so there are x*(x+1) seats total when you add up all the seats in the x buses

jimthompson5910 (jim_thompson5910):

that's what I was going for

jimthompson5910 (jim_thompson5910):

and x*(x+1) = x^2 + 1 is false it should be x*(x+1) = x^2 + x

jimthompson5910 (jim_thompson5910):

you have to distribute that x through to ALL of the terms inside

OpenStudy (anonymous):

Ooh ok. but I just put the 1 instead of x. my mistake @jim_thompson5910